OpenStudy (anonymous):

finde the derivative of x^2*(x-y)^2=X^2- Y^2

OpenStudy (anonymous):

wrt \(x\)?

OpenStudy (anonymous):

the right hand side is easiest, you get \(2x-2yy'\)

OpenStudy (anonymous):

the left hand side requires the product rule

OpenStudy (anonymous):

\[2x(x-y)^2+x^2\left(2(x-y)(1-y')\right)\]

OpenStudy (anonymous):

set \[2x(x-y)^2+x^2\left(2(x-y)(1-y')\right)=2x-2yy'\] and to a raft of algebra to solve for \(y'\)

OpenStudy (anonymous):

\[\frac{ -2x^3+3x^2y-xy^2+x }{ x^2y-x^3+y }\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

lol i have no idea

OpenStudy (anonymous):

i can't do all that algebra without making a mistake you can check your answer here http://www.wolframalpha.com/input/?i=x^2*%28x-y%29^2%3Dx^2-+y^2+ look at the very bottom for \(\frac{\delta y(x)}{\delta(x)}\)